# PCoreSet: Effective Active Learning through Knowledge Distillation from Vision-Language Models

### v1

@@ Abstract
Knowledge distillation (KD) is a widely used framework for training compact, task-specific models by leveraging the knowledge of teacher models. However, its application to active learning (AL), which aims to minimize annotation costs through iterative sample selection, remains underexplored. This gap stems from the fact that KD typically assumes access to sufficient labeled data, whereas AL operates in data-scarce scenarios where task-specific teacher models are often unavailable. In this paper, we introduce ActiveKD, a framework that integrates AL with KD by leveraging the zero- and few-shot capabilities of large vision-language models (VLMs). A key aspect of ActiveKD is the structured prediction bias of VLMs—i.e., their predictions form clusters in the probability space. We regard this structure as an inductive bias of the teacher model, capturing generalizable output patterns beneficial to student learning. To exploit this bias, we propose Probabilistic CoreSet (PCoreSet), a selection strategy that maximizes coverage in the probability space rather than the feature space. PCoreSet strategically selects categorically diverse unlabeled samples, facilitating more efficient transfer of teacher knowledge under limited annotation budgets. Evaluations on 11 datasets show that PCoreSet consistently outperforms existing selection methods within the ActiveKD framework, advancing research at the intersection of AL and KD.
@@ 1 Introduction
Recent advances in deep neural networks have focused on developing powerful generalist models capable of solving diverse tasks [2, 64, 55, 56] or creating robust transferable models through intensive pretraining [46, 67, 29, 17, 68, 38]. However, real-world applications often necessitate training compact task-specific models, with a significant challenge being task-specific data labeling due to high annotation costs in specialized domains. Active Learning (AL) addresses this by iteratively selecting the most informative samples for oracle annotation [71, 52], particularly in pool-based scenarios where informative samples can be identified from large unlabeled data. Semi-Supervised Learning (SSL) naturally complements AL by leveraging extensive unlabeled data during training [5, 10, 101, 41], enabling selection strategies tailored to the SSL framework [22, 53, 70, 79].
Figure 1: (Left): The proposed ActiveKD framework consistently improves the final-round accuracy of baseline selection methods, while PCoreSet outperforms other baselines when combined with ActiveKD. (Right): ActiveKD consistently improves the performance of PCoreSet across active learning rounds (No Distill vs. Zero-shot Distill), with further gains when using few-shot teachers (Few-shot Distill).
Knowledge distillation (KD) is a widely used framework for training compact models by transferring knowledge from large, pretrained teacher models [34]. Beyond its original purpose, KD naturally supports SSL by enabling the use of unlabeled data through teacher-generated soft labels [13, 32, 18, 94]. Despite its effectiveness, the application of KD in active learning (AL) remains underexplored. One major reason is the mismatch in assumptions: AL assumes that labels must be manually acquired for task-specific data through selective annotation, while KD typically assumes the existence of powerful task-specific teacher models trained on sufficient labeled data. However, in data-scarce settings common in AL, such teacher models are rarely available or practical to obtain.
@@
Recently, Vision-Language Models (VLMs) have emerged as powerful teachers in KD frameworks, as their rich representations—acquired through large-scale, unsupervised pretraining—enhance the transferability of knowledge across both modalities and tasks [84, 88, 59]. Recent work [41] has also shown distilling the knowledge of VLMs into compact, task-specific models effective in label-scarce scenarios, such as few-shot or low-shot semi-supervised settings. Motivated by these, we introduce ActiveKD, a framework that integrates AL with KD by leveraging the zero- and few-shot prediction capabilities of VLMs. Specifically, in each AL round, we train a task-specific student model using both the labeled data acquired up to that round and unlabeled data with soft predictions generated by the VLM teacher, as illustrated in Fig. 2.
@@
In ActiveKD settings, we identify an aspect of zero- or few-shot VLM teachers: they exhibit structured prediction biases acquired during pretraining and shaped by language prompts [6], resulting in predictions that form distinct clusters in the probability space. We observe that this bias propagates through distillation, influencing both the active selection mechanism and student model performance. Rather than viewing this as a limitation, we interpret this structure as an inductive bias of the teacher model—capturing generalizable patterns in the output space that, when complemented by labeled data, can benefit student learning. To exploit this, we propose a simple and effective method, Probabilistic CoreSet (PCoreSet), which selects coresets in the probability space rather than the feature space [76]. Specifically, it maximizes diversity in the categorical probability space by targeting underrepresented regions. In doing so, PCoreSet identifies samples that deviate from densely populated regions, effectively selecting underrepresented samples with limited budgets.
To evaluate the empirical effectiveness of ActiveKD and PCoreSet, we conduct extensive experiments on 11 datasets spanning a wide range of tasks, including generic object recognition [73, 20], fine-grained classification [63, 57, 65], domain-specific recognition [9], scene understanding [91], texture analysis [15], satellite imagery [33], and human action recognition [80]. As shown in Fig. 1, the results demonstrate that ActiveKD consistently improves the performance of active learning across all datasets, with further gains observed when using few-shot teachers. Moreover, the proposed PCoreSet outperforms baseline selection methods under ActiveKD, in both zero-shot and few-shot distillation settings. Our contributions and findings are summarized as follows:
@@
  • We introduce Active Knowledge Distillation (ActiveKD), a framework that leverages the zero- and few-shot capabilities of Vision-Language Models (VLMs) to integrate KD into Active Learning (AL), enabling efficient training of compact models with limited labeled data.
@@
  • We identify structured prediction biases in VLM teachers as an inductive bias that captures generalizable output patterns. We propose Probabilistic CoreSet (PCoreSet), which selects underrepresented samples in the probability space, enabling effective knowledge transfer.
  • Extensive experiments across 11 datasets, including ImageNet, validate the effectiveness of both ActiveKD and PCoreSet in various AL scenarios under the presence of VLM teachers.
@@ 2 Related Works
Recent advances in machine learning have focused on training large foundation models that either transfer pretrained knowledge to specific tasks [29, 17, 68, 38, 67, 46] or directly apply to target tasks [87, 55, 56, 68, 38] via task instructions or zero-shot prediction using language prompts. In this paper, we utilize VLMs [68, 38, 77] as teachers, which learn a shared embedding space for images and texts through a contrastive objective. A key advantage of such language-aligned training is the ability to perform zero-shot predictions [68, 38] without task-specific training and few-shot predictions [103, 102, 42, 104, 43, 100, 72, 98, 49, 23, 99, 96] with few examples, qualifying VLMs as generalist models capable of handling diverse visual recognition tasks. Despite their strong task-agnostic capabilities, effectively leveraging task-specific datasets remains crucial for developing compact models that perform well on downstream tasks, especially under label scarcity—a common challenge in real-world applications.
Active learning [AL; 71, 97, 52]
is a framework specifically designed to address such label-scarcity issues. It assumes that acquiring domain-specific datasets through manual annotation can be prohibitively expensive in practical scenarios. Specifically, AL addresses this by strategically selecting the most informative samples from unlabeled data on the target task for annotation, maximizing model performance with minimal labeled data. Previous approaches include uncertainty-based methods [51, 40, 35, 36, 21, 45, 69], diversity-based techniques [76, 66, 95, 27], and hybrid approaches [4, 45, 26, 24] that combine multiple criteria for sample selection. Several studies have focused on the class imbalance problem [3, 8, 37], where unbalanced datasets can lead selection algorithms to pick class-imbalanced samples. Recent research has integrated AL with prompt tuning [50, 39, 102] of VLMs, leveraging foundational knowledge with efficient parameter updates [6, 74, 44]. PCB [6] addressed skewed predictions of VLMs by balancing class distributions, while CB+SQ [44] enhanced this with class-guided clustering and adaptive thresholds. In contrast, we regard structured predictions as an inductive bias, and propose to leverage it by selecting probabilistically diverse samples rather than exploiting diversity in the feature space [76].
@@
Knowledge Distillation (KD) [34] transfers knowledge from large teachers to compact students. With the emergence of vision foundation models through self-supervised learning [12, 30, 25, 14, 11, 31] and vision-language pretraining [68, 38], researchers have increasingly explored KD methods to leverage knowledge embedded within these models. Early work focused on self-supervised models [19, 1, 92, 86, 62, 78], while the advent of VLMs inspired further research, including distillation of compact VLMs from larger counterparts [90, 81, 82, 83, 93], unsupervised distillation from VLM predictions [84, 88, 59], and few-shot semi-supervised distillation of VLMs [41]. Some research has explored using teacher models instead of human annotation to address incorrect predictions [7] or generate data efficiently [85, 54]. Our work focuses on human-in-the-loop scenarios, aiming to identify the most informative samples for annotation within a KD framework, by leveraging the zero- or few-shot capabilities of VLMs, to maximize learning efficiency under practical constraints of AL settings.
@@ 3.1 Preliminaries
We utilize Vision-Language Models (VLMs) like CLIP [68] and ALIGN [38] as teacher models. These models jointly optimize an image encoder $f_{\mathcal{X}}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ and a text encoder $f_{\mathcal{T}}:\mathcal{T}\rightarrow\mathbb{R}^{d}$ to map corresponding image-text pairs into a shared embedding space $\mathbb{R}^{d}$. This cross-modal alignment enables zero-shot transfer through natural language supervision. For $C$-way classification tasks, we create textual prompts (e.g., “a photo of a [CLASS]”) to generate class-specific text descriptions $\{t_{1},t_{2},\ldots,t_{C}\}$. The output probability vector of categorical distribution over $C$ classes is computed as the following:
@@
Here, $\Delta^{C-1}$ is the $(C{-}1)$-dimensional probability simplex, $\sigma:\mathbb{R}^{C}\to\Delta^{C-1}$ denotes the softmax function, and $\zeta\in\mathbb{R}_{>0}$ is the temperature scaling factor [34]. Cosine similarity is defined as $\mathtt{CosSim}(x,y)=\frac{x^{\top}y}{\|x\|_{2}\|y\|_{2}}$. The final predicted class is given by $\arg\max_{c\in[C]}[f(x)]_{c}$.
Figure 3: (Left): Teacher model prediction biases ($(\mu_{1},r_{1}),\ldots,(\mu_{k},r_{k})$) transfer to student models via distillation, where $\hat{\mu}_{k}=\mu_{k}+y_{c}$ and $\hat{r}_{k}=(1-\lambda)r_{k}$ ($\mu$ denotes centroids, $r$ denotes radii). (Right): PCoreSet selects samples maximizing distance to labeled points in probability simplex $\Delta^{C-1}$, uncovering underrepresented regions.
@@
1:Initial labeled dataset $\mathcal{D}^{(l)}$, unlabeled pool $\mathcal{D}^{(u)}$, teacher VLM $f$, an selection algorithm $A$, query size $Q$, number of rounds $R$
2:The final model $f_{r}$
@@
6:    Train $f_{r}$ with $\mathcal{L}_{\text{CE}}$ in Eq. 2, and $\mathcal{L}_{\text{KD}}$ in Eq. 3
@@
1:Labeled dataset $\mathcal{D}^{(l)}$, unlabeled pool $\mathcal{D}^{(u)}$, model $f_{r}$, query size $Q$, distance function $d(x,x^{\prime}):=||f_{r}(x)-f_{r}(x^{\prime})||_{2}$
2:Coreset $S\subset\mathcal{D}^{(u)}$ with $|S|=Q$
3:Initialize $\mathcal{S}=\mathcal{D}^{(l)}$
4:Initialize $D[x]=\min_{s\in\mathcal{S}}d(x,s),\forall x\in\mathcal{D}^{(u)}$
5:while $|\mathcal{S}|<Q$ do
6:    $x^{*}=\underset{x\in\mathcal{D}^{(u)}\setminus\mathcal{S}}{\arg\max}\,D[x]$
7:    $\mathcal{S}=\mathcal{S}\cup\{x^{*}\}$
@@
In this paper, we consider a pool-based active learning (AL), a framework for building a $C$-way classifier $f_{r}:\mathcal{X}\mapsto\Delta^{C-1}$ for each round $r\in[R]:=\{1,\ldots,R\}$ while minimizing annotation costs. We start with a labeled dataset $\mathcal{D}^{(l)}=\{(x_{n}^{(l)},y_{n})\}_{n=1}^{N}$ and an unlabeled dataset $\mathcal{D}^{(u)}=\{x_{m}^{(u)}\}_{m=1}^{M}$, where $y_{n}\in\{0,1\}^{C}$ is the one-hot encoding label of $x_{n}$ and typically $N\ll M$. We first train a model $f_{r}$ using the current labeled dataset $\mathcal{D}^{(l)}$ for each round $r$. Then, we select subset of unlabeled data that will be requested for annotation, i.e., $\{x_{q}^{*}\}_{q=1}^{Q}\leftarrow A(\mathcal{D}^{(l)},\mathcal{D}^{(u)},f_{r};Q)\subset\mathcal{D}^{(u)}$, where $Q$ is the number of query datapoints and $A$ is an algorithm for selection. These selected queries are then annotated by an oracle (typically human experts) to obtain their true labels $\{y_{q}^{*}\}_{q=1}^{Q}$. The labeled dataset is updated as $\mathcal{D}^{(l)}\leftarrow\mathcal{D}^{(l)}\cup\{(x_{q}^{*},y_{q}^{*})\}_{q=1}^{Q}$, and the unlabeled pool is reduced by $\mathcal{D}^{(u)}\leftarrow\mathcal{D}^{(u)}\setminus\{x_{q}^{*}\}_{q=1}^{Q}$. See Alg. 1 without underlines for an overview.
3.2 Our Method: ActiveKD with PCoreSet

### v2

# PCoreSet: Effective Active Learning through Knowledge Distillation from Vision-Langauge Models

@@ Abstract
Knowledge distillation (KD) is a widely used framework for training compact, task-specific models by transferring the knowledge from teacher models. However, its application to active learning (AL), which aims to minimize annotation costs through iterative sample selection, remains underexplored. This gap stems from the fact that KD typically assumes access to sufficient labeled data, whereas AL operates in data-scarce scenarios where task-specific teacher models are often unavailable. In this paper, we first introduce ActiveKD, a framework that integrates AL with KD by leveraging the zero- and few-shot capabilities of large vision-language models (VLMs). A key aspect of ActiveKD is the structured prediction bias of VLMs—i.e., their predictions form clusters in the probability space. We regard this structure as an inductive bias of the teacher model, capturing generalizable output patterns beneficial to student learning. To exploit this bias, we propose Probabilistic CoreSet (PCoreSet), a selection strategy that maximizes coverage in the probability space rather than the feature space. PCoreSet strategically selects probabilistically diverse unlabeled samples, facilitating more efficient transfer of teacher knowledge under limited annotation budgets. Extensive evaluations on 11 datasets show that ActiveKD consistently improves performance across selection methods (e.g., +29.07% on ImageNet, averaged over methods). Under ActiveKD, PCoreSet ranks first in 64/73 settings ($\approx$87.7%) across 5 student and 3 teacher networks, always achieving the best performance except for first 2 AL rounds. Our code is available at https://github.com/erjui/PCoreSet.
@@ 1 Introduction
Recent advances in deep neural networks have focused on developing powerful generalist models capable of solving diverse tasks (Achiam et al., 2023; OpenAI, 2025; Liu et al., 2023; 2024b) or creating robust transferable models through intensive pretraining (Koroteev, 2021; Radford et al., 2019; He et al., 2016; Dosovitskiy et al., 2020; Radford et al., 2021; Jia et al., 2021). However, real-world applications often necessitate training compact task-specific models, with a significant challenge of obtaining task-specific labeled data due to high annotation costs. Active learning (AL) addresses this by iteratively selecting the most informative samples for oracle annotation (Ren et al., 2021; Li et al., 2024), particularly in pool-based scenarios where informative samples can be identified from large unlabeled data. The availability of unlabeled data allows semi-supervised learning (SSL) to be applied to AL settings during training (Assran et al., 2021; Cai et al., 2022; Zheng et al., 2023; Kang et al., 2025), accompanying researches at the intersection of AL and SSL (Gao et al., 2020; Lim et al., 2023; Rangnekar et al., 2023; Singh et al., 2024).
Knowledge distillation (KD), meanwhile, is a widely used framework for training models by transferring knowledge from large pretrained teacher models (Hinton et al., 2015). Beyond its original purpose, KD can be naturally integrated into SSL framework as teacher models can provide its knowledge in the form of soft labels for unlabeled data (Chen et al., 2020b; He et al., 2021; Du et al., 2023; Yang et al., 2025). However, despite its effectiveness, the application of KD in AL remains underexplored. One major reason is the mismatch in their assumptions: AL assumes that labels must be manually acquired through selective annotation, while KD typically assumes the existence of powerful task-specific teacher models trained on sufficient labeled data. However, considering the data-scarce settings common in AL, such teacher models are rarely available or practical to obtain.
Figure 1: (Left): The proposed ActiveKD framework consistently improves the final-round accuracy of all selection methods, while PCoreSet further outperforms baselines when combined with ActiveKD. (Right): ActiveKD consistently improves the performance of PCoreSet across active learning rounds (No Distill vs. Zero-shot Distill), with further gains when using few-shot teachers (Few-shot Distill).
@@
Recently, vision-language models (VLMs) have emerged as powerful teachers in KD frameworks, as their rich representations—acquired through large-scale, unsupervised pretraining—enable effective knowledge transfer across both modalities and tasks (Vemulapalli et al., 2023; Wu et al., 2024; Mistretta et al., 2024). Recent work (Kang et al., 2025) has further demonstrated that transferring the knowledge of VLMs into task-specific models is effective in label-scarce scenarios, such as few-shot or low-shot semi-supervised settings. Motivated by these, we introduce ActiveKD, a framework that integrates AL with KD by leveraging the zero- and few-shot capabilities of VLMs. Specifically, in each AL round, we train a task-specific student model using both labeled data acquired up to that round and unlabeled data with soft predictions generated by the VLM teacher, as illustrated in Fig.˜2.
@@
In ActiveKD settings, we identify an aspect of VLM teachers: they exhibit structured prediction biases acquired during pretraining and shaped by language prompts (Bang et al., 2024), resulting in predictions that form distinct clusters in the probability space as illustrated in Fig.˜3. We observe that this bias propagates through distillation, influencing both the active selection mechanism and student model performance. Rather than viewing this as a limitation, we interpret this structure as an inductive bias of the teacher model—capturing generalizable patterns in the output space that, when complemented by labeled data, can benefit student learning. To exploit this, we propose a simple and effective method, Probabilistic CoreSet (PCoreSet), which selects coresets in the probability space rather than the feature space (Sener & Savarese, 2017). Specifically, it maximizes diversity in the categorical probability space by targeting underrepresented regions with limited budgets.
To evaluate the empirical effectiveness of ActiveKD and PCoreSet, we conduct extensive experiments on 11 datasets spanning diverse domains and tasks detailed in Appendix˜C. As summarized in Fig.˜1, ActiveKD consistently improves performance across selection methods on all datasets (e.g., +29.07% on ImageNet, averaged over methods), with further gains when using few-shot teachers. Moreover, under ActiveKD, PCoreSet outperforms alternative selection strategies, ranking first in 64/73 settings ($\approx$ 87.67%) across 5 student and 3 teacher architectures, with either 8 or 16 rounds.
Our contributions and findings are summarized as follows:
@@
  • We first introduce active knowledge distillation (ActiveKD), a framework that integrates knowledge distillation into active learning by leveraging the zero- and few-shot capabilities of vision-language models to train efficient task-specific models with limited labeled data.
@@
  • We identify structured prediction biases in VLM teachers that are propagated to the student model after KD. We thus propose Probabilistic CoreSet (PCoreSet), which utilizes this inductive bias to select samples in the underrepresented regions of the probability space.
  • Extensive experiments on 11 datasets demonstrate the effectiveness of the proposed ActiveKD and PCoreSet: ActiveKD consistently improves performance across all selection methods on all datasets (e.g., +29.07% on ImageNet, averaged over methods). PCoreSet also consistently outperforms alternative selection strategies under the ActiveKD framework, ranks first in 64/73 settings ($\approx$ 87.67%) across 5 student and 3 teacher architectures.
@@ 2 Related Works
Recent advances in machine learning have focused on training large foundation models that either transfer pretrained knowledge to specific tasks (He et al., 2016; Dosovitskiy et al., 2020; Radford et al., 2021; Jia et al., 2021; Radford et al., 2019; Koroteev, 2021) or directly apply to target tasks (Wei et al., 2021; Liu et al., 2023; 2024b; Radford et al., 2021; Jia et al., 2021) via task instructions or zero-shot prediction using language prompts. In this paper, we utilize VLMs (Radford et al., 2021; Jia et al., 2021; Silva-Rodriguez et al., 2024) as teachers, which learn a shared embedding space for images and texts through a contrastive objective. A key advantage of such language-aligned training is the ability to perform zero-shot predictions (Radford et al., 2021; Jia et al., 2021) without task-specific training and few-shot predictions (Zhou et al., 2022b; a; Khattak et al., 2023a; Zhu et al., 2023; Khattak et al., 2023b; Zhao et al., 2024; Roy & Etemad, 2023; Zhang et al., 2024; Lafon et al., 2025; Gao et al., 2024; Zhang et al., 2021; Yu et al., 2023) with few examples, qualifying VLMs as generalist models capable of handling diverse visual recognition tasks. Despite their strong task-agnostic capabilities, effectively leveraging task-specific datasets remains crucial for developing compact models that perform well on downstream tasks, especially under labeled data scarcity—a common challenge in real-world applications.
Active learning (AL; Ren et al., 2021; Zhan et al., 2022; Li et al., 2024) is a framework specifically designed to address such labeled data scarcity issues. It assumes that acquiring domain-specific datasets through manual annotation can be prohibitively expensive in practical scenarios. Specifically, AL addresses this by strategically selecting the most informative samples from unlabeled data on the target task for annotation, maximizing model performance with minimal labeled data. Previous approaches include uncertainty-based methods (Lewis, 1995; Joshi et al., 2009; Holub et al., 2008; Houlsby et al., 2011; Gal et al., 2017; Kirsch et al., 2019; Rakesh & Jain, 2021), diversity-based techniques (Sener & Savarese, 2017; Parvaneh et al., 2022; Yehuda et al., 2022; Hacohen et al., 2022), and hybrid approaches (Ash et al., 2019; Kirsch et al., 2019; Hacohen & Weinshall, 2023; Giouroukis et al., 2025) that combine multiple criteria for sample selection. Several studies have focused on the class imbalance problem (Aggarwal et al., 2020; Bengar et al., 2022; Huang et al., 2024), where unbalanced datasets can lead selection algorithms to pick class-imbalanced samples. Recent research has integrated AL with prompt tuning (Lester et al., 2021; Jia et al., 2022; Zhou et al., 2022a) of VLMs, leveraging foundational knowledge with efficient parameter updates (Bang et al., 2024; Safaei & Patel, 2024; Kim et al., 2024). PCB (Bang et al., 2024) addressed skewed predictions of VLMs by balancing class distributions, while CB+SQ (Kim et al., 2024) enhanced this with class-guided clustering and adaptive thresholds. In contrast, we regard structured predictions as an inductive bias, and propose to leverage it to select samples to annotate.
@@
Knowledge Distillation (KD) (Hinton et al., 2015) transfers knowledge from large teachers to compact students. With the emergence of vision foundation models through self-supervised learning (Chen et al., 2020a; He et al., 2020; Grill et al., 2020; Chen & He, 2021; Caron et al., 2021; He et al., 2022) and vision-language pretraining (Radford et al., 2021; Jia et al., 2021), researchers have explored KD methods to leverage knowledge embedded within these models, moving beyond conventional KD that require training student models on identical datasets with substantial data. Early work focused on self-supervised models (Fang et al., 2021; Abbasi Koohpayegani et al., 2020; Xu et al., 2021; Wang et al., 2022; Navaneet et al., 2022; Singh & Wang, 2024), while the advent of VLMs inspired further research, including distillation of compact VLMs from larger counterparts (Wu et al., 2023; Sun et al., 2023; Udandarao et al., 2024; Vasu et al., 2024; Yang et al., 2024), unsupervised distillation from VLM predictions (Vemulapalli et al., 2023; Wu et al., 2024; Mistretta et al., 2024), and few-shot semi-supervised distillation of VLMs (Kang et al., 2025). Some research has explored using teacher models instead of human annotation to address incorrect predictions (Baykal et al., 2022) or generate data efficiently (Wang et al., 2020; Liu et al., 2024a). Our work focuses on human-in-the-loop scenarios, aiming to identify the most informative samples for annotation within a KD framework, by leveraging the zero- or few-shot capabilities of VLMs, to maximize learning efficiency under practical constraints of AL settings.
@@ 3.1 Preliminaries
We utilize Vision-Language Models (VLMs) like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) as teacher models. These models jointly optimize an image encoder $f_{\mathcal{X}}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ and a text encoder $f_{\mathcal{T}}:\mathcal{T}\rightarrow\mathbb{R}^{d}$ to map corresponding image-text pairs into a shared embedding space $\mathbb{R}^{d}$. This cross-modal alignment enables zero-shot transfer through natural language supervision. For $C$-way classification tasks, we create textual prompts (e.g., “a photo of a [CLASS]”) to generate class-specific text descriptions $\{t_{1},t_{2},\ldots,t_{C}\}$. The output probability vector of categorical distribution over $C$ classes is computed as the following:
@@
Here, $\Delta^{C-1}$ is the (C1)(C{-}1)-dimensional probability simplex, $\sigma:\mathbb{R}^{C}\to\Delta^{C-1}$ denotes the softmax function, and $\zeta\in\mathbb{R}_{>0}$ is the temperature scaling factor (Hinton et al., 2015). Cosine similarity is defined as $\mathtt{CosSim}(x,y)=\frac{x^{\top}y}{\|x\|_{2}\|y\|_{2}}$. The final predicted class is given by $\arg\max_{c\in\{1,\ldots,C\}}[f(x)]_{c}$.
@@
1:Initial labeled dataset $\mathcal{D}^{(l)}$, unlabeled pool $\mathcal{D}^{(u)}$, teacher VLM $f$, a selection algorithm AA, query size $Q$, number of rounds $R$
2:Final model $f_{r}$
@@
6:  Train $f_{r}$ with $\mathcal{L}_{\text{CE}}$ in Eq.˜2 and $\mathcal{L}_{\text{KD}}$ in Eq.˜3
@@
1:Labeled dataset $\mathcal{D}^{(l)}$, unlabeled pool $\mathcal{D}^{(u)}$, model $f_{r}$, query size $Q$, distance $d(x,x^{\prime}):=\lVert f_{r}(x)-f_{r}(x^{\prime})\rVert_{2}$
2:Coreset $\mathcal{S}\subset\mathcal{D}^{(u)}$ with $\lvert\mathcal{S}\rvert=Q$
3:Initialize $\mathcal{S}\leftarrow\mathcal{D}^{(l)}$
4:Initialize $D[x]\leftarrow\min_{s\in\mathcal{S}}d(x,s)$ for all $x\in\mathcal{D}^{(u)}$
5:while $\lvert\mathcal{S}\rvert<Q$ do
6:  $x^{*}\leftarrow\arg\max_{x\in\mathcal{D}^{(u)}\setminus\mathcal{S}}D[x]$
7:  $\mathcal{S}\leftarrow\mathcal{S}\cup\{x^{*}\}$
@@
Figure 3: (Left): Teacher model prediction biases ($(\mu_{1},r_{1}),\ldots,(\mu_{k},r_{k})$) are transferred to student models via distillation, where $\hat{\mu}_{k}=\mu_{k}+y_{c}$ and $\hat{r}_{k}=(1-\lambda)r_{k}$ ($\mu$ denotes centroids, $r$ denotes radii). (Right): PCoreSet selects samples maximizing distance to labeled points in probability simplex $\Delta^{C-1}$, uncovering underrepresented regions.
@@
In this paper, we consider a pool-based active learning (AL), a framework for building a $C$-way classifier $f_{r}:\mathcal{X}\mapsto\Delta^{C-1}$ for each round $r\in\{1,\ldots,R\}$ while minimizing annotation costs. We start with a labeled dataset $\mathcal{D}^{(l)}=\{(x_{n}^{(l)},y_{n})\}_{n=1}^{N}$ and an unlabeled dataset $\mathcal{D}^{(u)}=\{x_{m}^{(u)}\}_{m=1}^{M}$, where $y_{n}\in\{0,1\}^{C}$ is the one-hot encoding label of $x_{n}$ and typically $N\ll M$. We first train a model $f_{r}$ using the current labeled dataset $\mathcal{D}^{(l)}$ for each round $r$. Then, we select subset of unlabeled data that will be requested for annotation, i.e., $\{x_{q}^{*}\}_{q=1}^{Q}\leftarrow A(\mathcal{D}^{(l)},\mathcal{D}^{(u)},f_{r};Q)\subset\mathcal{D}^{(u)}$, where $Q$ is the number of query datapoints and $A$ is an algorithm for selection. These selected queries are then annotated by an oracle (typically human experts) to obtain their true labels $\{y_{q}^{*}\}_{q=1}^{Q}$. The labeled dataset is updated as $\mathcal{D}^{(l)}\leftarrow\mathcal{D}^{(l)}\cup\{(x_{q}^{*},y_{q}^{*})\}_{q=1}^{Q}$, and the unlabeled pool is reduced by $\mathcal{D}^{(u)}\leftarrow\mathcal{D}^{(u)}\setminus\{x_{q}^{*}\}_{q=1}^{Q}$. See Algorithm˜1 without underlines for an overview.
3.2 ActiveKD